# Contextualizing Recommendation Explanations with LLMs: A User Study

### v1

@@ 3.6 Ethical Considerations
We respect the privacy and agency of all subjects potentially impacted by this work and take steps to protect their privacy. The experiment design has obtained approval from the Ethics Commission at the University of Lausanne. All data were collected anonymously, and any personally identifiable features were removed. All steps during data collection and analyses followed standards for ethical research (Rivers and Lewis 2014). We also informed users that movie recommendations and explanations were generated by an LLM.
@@ 4.1 Cognitive Needs
We observed a significant difference (Holm-adjusted $p<0.05$) in the distribution of the aggregated Cognitive measure (i.e., Scrutability, Transparency, Trust) between NE ($\mu=4.87$, $95\%$ CI = [4.78, 4.95]) and CE ($\mu=5.22$, $95\%$ CI = [5.14, 5.30]). We also found a significant difference between GE and CE (Holm-adjusted $p<0.05$), with CE surpassing GE ($\mu=4.90$, $95\%$ CI = [4.81, 4.99]). However, there was no statistical difference between NE and GE (Holm-adjusted $p>0.05$), suggesting that generic explanations do not differ markedly from no explanations in meeting users’ cognitive needs.
In the mixed-effects regression, both RecSysUtility ($\beta=0.319$, $p<0.05$) and MovieConsumption ($\beta=0.149$, $p<0.05$) demonstrated positive, significant effects on Cognitive, indicating that participants who perceive the recommender system as more useful and who watch movies more frequently tend to rate higher in Cognitive.
In the follow-up interviews, several participants expressed increased trust and clarity when receiving the contextualized explanations. One participant noted: “When it told me the rationale for suggesting the movie, I could decide whether or not to watch it.” Another participant stated: “Including the movie I had watched and comparing the similarities made me trust the recommendation more.” Interestingly, some users also noted that references to movies they disliked helped them filter out unwanted content, improving their sense of control and overall transparency.
These quantitative and qualitative findings suggest that contextualized explanations, by explicitly linking a user’s past and present preferences, can bolster trust and understanding in a way that generic or absent explanations do not. The more personal and concrete presentation of information serves users’ cognitive needs by explaining why some recommendations might be a good match, and others might not align with their tastes, ultimately leading to a deeper sense of scrutiny, transparency, and trust.
Figure 2: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
@@ 4.2 Affective Needs
We found no overall difference in the aggregated Affective measure (i.e., Persuasiveness, Satisfaction) among NE ($\mu=4.97$, $95\%$ CI = [4.88, 5.05]), GE ($\mu=4.92$, $95\%$ CI = [4.83, 5.01]), and CE ($\mu=5.02$, $95\%$ CI = [4.93, 5.11]). However, examining each subdimension showed that Persuasiveness differed significantly between GE and CE (Holm-adjusted $p<0.05$), indicating that contextualized explanations can make recommendations more compelling than generic explanations, even if this advantage does not fully carry over to overall perceived affective needs.
The mixed-effects regression highlights two user-engagement variables, RecSysUtility ($\beta=0.311$, $p<0.05$) and MovieConsumption ($\beta=0.167$, $p<0.05$), as positive, significant predictors of affective responses. This suggests that users who view the recommender system as helpful and frequently watch movies are more likely to feel emotionally drawn in, regardless of the explanation style.
Our interviews offer deeper insights into how these affective needs take shape. Several participants emphasized that a good explanation heightened their eagerness to watch the movie: “They made me want to watch them more than I would have before with their explanations.” Others found “overly pushy” or “over the top” descriptions off-putting, highlighting the delicate balance between building excitement and overselling the film. Some appreciated specific references to actors or plot details, while others disliked spoiler-like content, favoring a brief, suggestive teaser. These varying reactions prove that users’ emotional engagement hinges on contextual cues and personal taste, mood, and openness to promotional language.
These findings imply that satisfying affective needs, such as feeling entertained, inspired, or emotionally “in tune”, goes beyond simply adding contextual detail. Users’ enjoyment levels are tightly bound to their existing trust in the system, general interest in watching movies, and thresholds for promotional or spoiler content. Consequently, while contextualized explanations enhance persuasiveness for some users, effectively catering to affective motivations requires a more nuanced approach that recognizes diverse emotional triggers and preferences.
Figure 3: Affective ratings by explanation type (shown variable: Affective).
@@ 4.3 Utilitarian Needs
There is no overall difference in the aggregated Utilitarian measure (i.e., Effectiveness, Efficiency, Helpfulness) across NE ($\mu=5.24$, $95\%$ CI = [5.16, 5.32]), GE ($\mu=5.27$, $95\%$ CI = [5.19, 5.36]), and CE ($\mu=5.27$, $95\%$ CI = [5.18, 5.35]). However, a closer examination at Helpfulness revealed that NE ($\mu=5.07$, $95\%$ CI = [4.98, 5.17]) versus CE ($\mu=5.23$, $95\%$ CI = [5.13, 5.33]) was significant (Holm-adjusted $p<0.05$), suggesting that contextualized explanations can offer an advantage over having no explanation in practice by helping users assess whether a movie meets their needs. By contrast, there was no difference between GE and CE in Helpfulness, indicating that once any form of explanation is provided, adding contextual details may not further enhance users’ perceived practical value in selecting movies.
The mixed-effects regression showed RecSysUtility ($\beta=0.292,p<0.05$) and MovieConsumption ($\beta=0.181,p<0.05$) both positively influenced Utilitarian, again implying that users who closely engage with recommender systems and movies, tend to see greater practical value in the recommendations regardless of explanation style. Female participants also reported slightly higher utilitarian scores (versus male, $p<0.05$).
Our interview provides further evidence. Several participants appreciated having clear, concise details about the plot or genre, emphasizing how these insights save time in deciding whether a movie fits their tastes: “The more it detailed the movie, the more it helped me decide.” Some also suggested that referencing both liked and disliked movies aided them in filtering out poor matches and reinforcing good ones: “I was able to assess whether I wanted to watch it, based on what I liked before, and also on what I knew I didn’t enjoy.” Others felt the explanations were too vague or enthusiastic, hindering their ability to gauge true quality.
Although contextualized explanations may slightly outperform no explanations in offering helpful guidance, overall utilitarian gains do not appear large when considering Effectiveness or Efficiency. These findings suggest that users’ utilitarian needs stem less from how the explanation is framed and more from whether the system is already perceived as useful and aligns with existing consumption habits. Nonetheless, highlighting specific points of similarity or dissimilarity can provide an additional benefit in helping users quickly identify which recommendations are worth pursuing and which to skip.
Figure 4: Utilitarian ratings by explanation type (shown variable: Utilitarian).
@@ 4.4 Consumption Intentions
Knowing users’ consumption intentions (measured by UserIntent) helps us understand how explanations translate into actual viewing decisions. Pairwise comparisons confirm that CE ($\mu=4.70$, $95\%$ CI = [4.59, 4.81]) significantly differs from GE ($\mu=4.31$, $95\%$ CI = [4.18, 4.43]) (Holm-adjusted $p<0.05$). The contrasts between NE ($\mu=4.53$, $95\%$ CI = [4.42, 4.65]) and CE, as well as NE and GE, are also significant (Holm-adjusted $p<0.05$). This pattern suggests that the existence of explanations and contextualization fosters greater motivation to watch a recommended movie.
Mixed-effects regression indicates that CE increases consumption intentions by 0.240 points on average ($\beta=0.240,p<0.05$) compared to GE. RecSysUtility ($\beta=0.336,p<0.05$) and MovieConsumption ($\beta=0.143,p<0.05$) both increase the likelihood of watching the recommended title. Meanwhile, older participants ($\beta=-0.139,p<0.05$) and female users ($\beta=-0.227,p<0.05$) expressed slightly lower consumption intentions.
The qualitative interview responses underscore how alignment with personal taste can motivate users to watch the recommended movies. Several participants mentioned that recognizable themes, plot twists, or well-known actors greatly influenced their choices. One participant explained: “An interesting twist or variation from an expected plot development intrigues me, as does the presence of an actor I like.” Meanwhile, others highlighted that the mention of disliked or irrelevant elements quickly led them to dismiss a recommendation: “Some movies I simply will not watch … if it’s not my kind of genre, no explanation can convince me.” In some instances, participants weighed social factors such as the opinions of friends and family members, underscoring how shared viewing decisions can override individual preferences if, for example, a partner or group prefers a different type of film.
These results imply that while contextualized explanations provide an edge over generic ones, individual beliefs about system helpfulness and prior consumption habits remain critical in determining whether users actually watch a recommended movie. Explaining “why this movie” helps, but so does the user’s fundamental perception of the recommender’s reliability and personal engagement with the medium. Ultimately, consumption intentions emerge from a synthesis of explanation quality, personal taste, and social or experiential factors, illustrating that users weigh both the persuasive power of contextualization and their unique circumstances.
Figure 5: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
@@
0.049
@@
--0.004
--0.043
0.106
MovieConsumption
0.149***
0.167***
0.181***
0.143*
DisplaypOrder
--0.018
--0.046
--0.031
--0.022
Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
Table 4: Mixed-effect linear regression results. Coefficients are shown with standard errors in parentheses.
5 Supplementary Experiments
To supplement the main experiment, we expanded the experiment to a larger scope by leveraging another LLM (Claude 3.5 Sonnet).
Given that GPT-4 Turbo is one of the most powerful LLMs, we tested whether the superior performance of contextualization could be generalized to other LLMs. We generated another recommendation corpus using Claude 3.5 Sonnet and replicated the experiments on the U.S. samples ($N=140$, 71 are women, 68 are men, and one is non-binary). The results are aligned with our main findings, validating the advantages of LLMs in generating contextualized explanations in terms of meeting users’ motivational needs and increasing their consumption intent. The results and interpretations are provided in Appendix C.
@@ 6.1 Implications of Findings
Recommender systems play a central role in shaping user decisions, particularly influential in the context of web and social media. Given the far-reaching implications of these systems, there is growing demand from users and regulators for transparency, emotional connection, or efficiency in decision-making. This highlights the value of high-quality and context-aware recommendation explanations. However, many existing recommender systems do not or fail to provide explanations that address users’ diverse needs (Li, Zhang, and Chen 2020; Lu et al. 2023).
@@
Our study demonstrates that LLM-generated contextualized explanations can overcome these limitations by satisfying users’ cognitive needs, ultimately increasing their consumption intentions. Also, our findings reveal the insufficiency of LLM-generated explanations in meeting users’ utilitarian and affective needs, which raises concerns about the proper design and implications of LLM-generated personalized content.
Benefits of contextualization. Our results show that contextualized explanations significantly satisfy users’ needs and their intention to watch the recommended movies. This is particularly evident in users’ cognitive needs (Scrutability, Transparency, Trust), where contextualization not only promotes a clearer understanding of why specific titles are recommended but also encourages users to scrutinize recommendations more carefully. The increase in trust and perceived transparency shows the importance of tailoring explanations to individual needs. Offering personally relevant and meaningful recommendations can encourage users to act on these suggestions, especially in entertainment and e-commerce settings, where a personalized touch can be key to customer retention.
LLMs in the recommendation. As LLMs evolve, their capacity to generate richer, more personalized explanations will continue to grow (Feuerriegel et al. 2025). This capability is particularly relevant for companies like Amazon and Netflix, where personalization is a competitive differentiator (Burroughs 2019). Our findings highlight how integrating LLMs into recommender systems can satisfy users’ cognitive needs and increase their consumption intention. Beyond driving engagement, context-aware explanations may help companies fulfill emerging regulatory requirements, such as the EU’s AI Act, which calls for transparency and the “right to explanation” in AI-driven decisions (Veale and Borgesius 2021; You and Suh 2024).
System utility and positive feedback loops. Our study suggests that a user’s active and positive engagement with the recommender system amplifies the benefits of contextualization. When participants view the system as helpful and relevant to their interests, context-aware explanations further boost their ratings of recommendation quality, trust, and helpfulness. This synergy can create a positive feedback loop: as users find the system increasingly beneficial, they engage more deeply, and contextualized explanations, in turn, reinforce these favorable perceptions. Moreover, frequent movie watchers also exhibit higher cognitive, affective, and utilitarian evaluations. Increasing user engagement through personalized interfaces, easy-to-use features, or curated content could, therefore, magnify the impact of contextualized explanations (Lacker and Way 2024). Companies seeking to cultivate sustained user relationships can harness this dynamic by combining context-rich recommendation details with user education or incentives that further encourage exploration and use of the system.
Avoiding redundant information. An interesting insight from the experiment is that some participants found overly “pushy” or “over the top” descriptions off-putting. In contrast, others appreciated detailed explanations. Although contextualized explanations can meet various user needs, avoiding irrelevant or excessive information that increases cognitive load is critical. Our findings suggest that, for specific users, no explanation is preferable to offering a generic one. This reflects a balancing act: well-crafted, context-specific explanations can significantly improve user perceptions, but lengthy or poorly targeted explanations can become a distraction.
In industries such as travel and finance, where decisions can be complex, selective use of contextualized explanations could help users make informed choices without overwhelming them. By highlighting only the most pertinent details, companies can streamline the decision-making process, thus promoting higher engagement and satisfaction.

### v2

@@ 3.6 Ethical Considerations
We respect the privacy and agency of all subjects potentially impacted by this work and take steps to protect their privacy. The experiment design has obtained approval from the Ethics Commission at the University of (name anonymized for review). All data were collected anonymously, and any personally identifiable features were removed. All steps during data collection and analyses followed standards for ethical research (Rivers and Lewis 2014). We also informed users that movie recommendations and explanations were generated by an LLM.
@@ 4 Results
Figure 2: Ratings by explanation type (shown variable: Cognitive, Affective, Utilitarian, and UserIntent). Significance levels between conditions are: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
@@ 4.1 Cognitive Needs
We observed a significant difference (Holm-adjusted $p<0.05$) in the distribution of the aggregated Cognitive measure between NE ($\mu$=4.94), GE ($\mu$=4.97), and CE ($\mu$=5.38). Pairwise tests showed that CE significantly outperformed both NE and GE (Holm-adjusted pp<0.05), indicating that contextual explanations better support users’ cognitive understanding. However, there was no statistical difference between NE and GE (Holm-adjusted pp>0.05), suggesting that generic explanations offer no additional cognitive benefit over providing no explanation.
In the mixed-effects regression, both RecSysUtility ($\beta$=0.304, pp<0.05) and MovieConsumption ($\beta$=0.150, pp<0.05) demonstrated positive, significant effects on Cognitive, indicating that participants who perceive the recommender system as more useful and who watch movies more frequently tend to rate higher in Cognitive. Besides, more movie contexts ($\beta$=0.048, pp<0.05) in the explanation also better satisfy users’ cognitive needs.
In the interviews, several participants clearly expressed their trust and commended the clarity when receiving contextualized explanations.
“The thing that makes me trust the explanation was reading how the description had brought up other titles of movies that I saw in the recommendations, that had true things about the characters, and how the movies could be similar or how I could watch them already.” (Female, 29)
“They made the details of the movies sound more exciting and interesting and gave you more of a full rundown of theories so you don’t think it’s going to be one thing and turn out another.” (Female, 31)
On the other hand, sometimes users disliked the recommendations, but they still helped to filter out unwanted content, improving their sense of control and overall transparency.
“The movies it recommended didn’t interest me, so I don’t think I’ll be watching them. Felt like I understood the movies weren’t for me.” (Male, 32)
“I trusted the recommended movie, but I made my own mind up after reading the explanation.” (Male, 44)
@@ 4.2 Affective Needs
For the Affective dimension, CE ($\mu$=5.19) was rated significantly higher than both NE ($\mu$=5.05) and GE ($\mu$=4.99) (Holm-adjusted pp<0.05). Still, no significant difference emerged between NE and GE (Holm-adjusted pp>0.05).
The mixed-effects regression highlights two user-engagement variables, RecSysUtility ($\beta$=0.295, pp<0.05) and MovieConsumption ($\beta$=0.175, pp<0.05), as positive, significant predictors of affective responses. This suggests that, regardless of the explanation style, users who view the recommender system as helpful and frequently watch movies are more likely to have positive feedback in affective needs. Besides, more movie context ($\beta$=0.069, pp<0.05) in the explanation has a positive effect as well.
From the interview, we observed that the movie and explanation quality are important to users’ affection, no matter whether it contains the context. This also explains why users’ satisfaction is fairly similar when receiving CE or GE. Regarding the textual quality, the output of LLMs is consistent regardless of the explanation types.
“Because I am a movie buff, and frankly, the descriptions were very well laid out so that it gave me enough information about the plot but also left enough unsaid that now I have to find out.” (Male, 39)
“Some of the word descriptions were a little over the top. The description was a little graphic.” (Female, 68)
“The explanation was well written, so it gave away a little more than I would normally like, but was just as good as a trailer in my mind.” (Male, 39)
Figure 3: Radar chart of average user perception ratings in different dimensions by explanation types.
@@ 4.3 Utilitarian Needs
In the Utilitarian dimension, CE ($\mu$=5.44) was rated significantly higher than NE ($\mu$=5.32) (Holm-adjusted pp<0.05). There was no significant difference between NE and GE or CE and GE ($\mu$=5.34).
The mixed-effects regression showed RecSysUtility ($\beta$=0.271, pp<0.05) and MovieConsumption ($\beta$=0.172, pp<0.05) both positively influenced Utilitarian, implying that users who closely engage with recommender systems and movies tend to see greater practical value in the recommendations regardless of explanation style. Female participants also reported higher utilitarian scores (versus male, $\beta$=0.292, pp<0.05). More movie context increases utilitarian value ($\beta$=0.044, pp<0.05).
In general, users felt that these explanations save time in deciding whether a movie fits their tastes. However, we observed that some users appreciated more details to support their decisions, while some preferred the explanations to be simplified.
“If I couldn’t clearly understand the explanation of the film, I thought that I was most likely not going to enjoy the movie. I trust my first instincts on things because it has saved me a lot of time in the past, by allowing me to avoid seeing a movie that would only waste time in my too short life.” (Female, 37)
“Sometimes too much is overwhelming.” (Male, 37)
“The more detailed movie, the more it helped me decide.”(87, Female)
@@ 4.4 Consumption Intentions
We found significant differences in the consumption intention across all conditions. CE ($\mu$=4.85) outperformed GE ($\mu$=4.36) and NE ($\mu$=4.61) (Holm-adjusted pp<0.05). GE was also rated significantly lower than NE (Holm-adjusted pp<0.05).
Mixed-effects regression indicates that RecSysUtility ($\beta$=0.292, pp<0.05) and MovieConsumption ($\beta$=0.125, pp<0.05) both increase the likelihood of watching the recommended title. Older participants ($\beta$=-0.147, pp<0.05) expressed lower consumption intentions. Besides, more movie context ($\beta$=0.101, pp<0.05) in the explanation has a positive effect as well.
Though users have overall positive feelings, their consumption intent is more attached to their taste. Many interviewees explicitly expressed that, if the movie is not their type, they will not watch it. This is not a consequence of our findings but of the accuracy of the algorithm behind the recommender systems. Hence, this still confirms the value of contextualization in motivating users’ decisions.
“The factor that motivated me the most in deciding whether or not to watch the movie was the correlation factor between the new movie and movies I have watched in the past.”(Female, 20)
“Sometimes the explanation was so intriguing I felt as though I wanted to know more about the movie. If the explanation is so boring, I don’t finish it, I’m most likely not going to watch the movie.” (Female, 37)
“If it seemed like the movie was for children, it was a no for me.”(Male, 32)
“This is becoming a little redundant, but again it explained it to the point I could see in my head, the picture and the action happening. Plus, anything with Will Smith has got to be awesome.” (Female, 61)
@@
0.038
0.028
0.003
0.040
RecSysUtility
0.304***
0.295***
0.271***
0.292***
RecSysFrequency
0.019
-0.004
-0.023
0.132
MovieConsumption
0.150**
0.175***
0.172***
0.125*
DisplayOrder
0.040
0.035
0.017
0.115**
HistoryNum
0.048**
0.069***
0.044**
0.101***
HistoryNum $\times$ MovieConsumption
-0.028
0.007
-0.032*
@@
-0.054
Table 4: Mixed-effects linear regression results. Coefficients are reported with standard errors in parentheses. Significance levels: $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
4.5 Interpretations
Based on the results above, we observed an overall advantage of LLM-generated contextualized explanations in addressing users’ motivational needs and increasing their consumption intentions. Figures 2 and 3 show that contextualized explanations successfully address users’ diverse needs. However, a key limitation remains: the explanations do not always align with users’ literary and genre preferences. Although we asked participants to evaluate the explanations rather than the recommended movies, interviews revealed that many users still evaluated the two together, resulting in lower scores for persuasiveness and consumption intention. When users turn to recommender systems, they often have specific goals, such as finding a movie to watch or exploring new options. Once the system delivers relevant recommendations, users may feel that their immediate needs are already satisfied. This likely explains why participants rated effectiveness, efficiency, and satisfaction high, even in the absence of explanations. Together, our findings highlight the value of LLM-generated explanations in building transparency and trust. However, to truly enhance user experience, explanations need to resonate with users’ tastes and expectations, especially in subjective domains like entertainment.
Demographics and prior engagement shape user perception: Based on the results in Table 4, users’ prior engagement significantly shapes their evaluation of explanations. Participants who consumed more movies rated explanations more positively across all dimensions, suggesting that domain familiarity enhances appreciation of contextualized content. Similarly, higher perceived usefulness of recommender systems strongly gave favorable evaluations.
Demographic effects were more selective. Female users gave higher utilitarian ratings, while older users rated affective and intention outcomes slightly lower. These findings suggest that both system engagement and user characteristics should be considered when designing a user-friendly system.
More viewing history contexts, better perceptions: Table 4 reveals that the number of viewing histroy (HistoryNum) positively influenced users’ evaluations in perception and intention. As detailed in Section 3.3, we limited the number of contexts maximum of four, which was guided by typical movie description lengths, native reading speeds, and cognitive load thresholds.
Importantly, the non-significant interaction effects between HistoryNum and both MovieConsumption and RecSysUtility suggest that the positive impact of including more movie contexts in explanations holds consistently, regardless of users’ movie consumption frequency or perceptions of the recommender system’s helpfulness. This implies that richer contextualization can enhance user evaluations without being contingent on prior engagement. Although richer context tends to improve perception and intent, the effects, based on the coefficients in the regression results, remain relatively small. Some users also flagged certain explanations as redundant or overly detailed. To prevent cognitive overload, we should carefully calibrate how much context to include, ensuring that explanations remain concise yet informative.
5 Heterogeneity Analysis
5.1 Claude 3.5 Sonnet
To supplement the main experiment, we expanded the experiment to a larger scope by leveraging another LLM (Claude 3.5 Sonnet). Given that GPT-4 Turbo is one of the most powerful LLMs, we tested whether the superior performance of contextualization could be generalized to other LLMs. We generated another recommendation corpus using Claude 3.5 Sonnet and replicated the experiments on the U.S. samples (NN=140, 71 are women, 68 are men, and one is non-binary). The results are aligned with our main findings, validating the advantages of LLMs in generating contextualized explanations in terms of meeting users’ motivational needs and increasing their consumption intent. The results and interpretations are provided in Appendix C.1.
5.2 Movie Genre
To further understand how user perceptions vary across recommendation contexts, we examined the role of movie genre, specifically comedy and adventure, with supplementary mixed-effects regression models (see Equation C.2 in Appendix C.2). As shown in Table 7, comedy movies were consistently rated lower than adventure movies across affective, utilitarian, and consumption intention outcomes. We also observed that these genre effects were not moderated by explanation types. These supplementary results highlight genre as a key factor influencing user responses to LLM-generated recommendations. The influence of movie genre remains an important area for further investigation.
@@ 6.1 Implications of Findings
Recommender systems play an important role in shaping user decisions, particularly influential in the context of web and social media. Given the far-reaching implications of these systems, there is growing demand from users and regulators for transparency, emotional connection, and efficiency in decision-making. This highlights the value of high-quality and context-aware recommendation explanations. However, many existing recommender systems do not or fail to provide explanations that address users’ diverse needs (Li, Zhang, and Chen 2020; Lu et al. 2023).
@@
Our study demonstrates that LLM-generated contextualized explanations can overcome these limitations by satisfying users’ cognitive needs, ultimately increasing their consumption intentions. Also, our findings reveal the insufficiency of LLM-generated explanations in meeting users’ utilitarian and affective needs, which raises concerns about the proper design and implications of LLM-generated contextualized content.
Build trust, but risk misrepresenting preferences: Our results show that contextualized explanations significantly address users’ cognitive needs, particularly in terms of Transparency, Trust, and Scrutability. By referencing users’ prior viewing history, LLM-generated explanations appear more tailored and meaningful, encouraging deeper engagement with recommendations.
Nevertheless, we also observed that LLMs tend to interpret viewing histories as an implicit preference signal. While this can increase perceived personalization, it also introduces a risk of over-interpretation, which is reflected in user interviews. In practice, real-world recommender systems often infer preferences from behavioral signals (e.g., clicks, watch time) rather than explicit feedback, which may not always align with users’ true interests or sentiments. LLMs, when generating explanations based on such data, may reinforce these weak or ambiguous signals, thereby creating a disconnect between the explanation and the user’s actual intent. This misalignment risks undermining the trust that contextualized explanations build. When users feel the system attributes preferences they do not hold, trust may erode over time. Therefore, in real-world deployments, systems that leverage LLMs for explanation generation should account for this interpretive bias and consider more nuanced strategies for referencing user history, particularly when preferences are inferred rather than explicitly stated.
Opportunities and challenges for the movie industry and users: Our findings show that LLM-generated movie recommendations significantly increase users’ intention to engage with suggested content. Effective recommendations can be generated by leveraging typical user viewing histories and state-of-the-art LLMs—without requiring additional user input or behavioral interventions. Such ease of implementation, coupled with LLMs’ ability to produce highly contextualized content, makes them valuable tools for streaming services and other media platforms.
Integrating LLMs into recommendation systems offers the potential to enhance operational efficiency but introduces significant costs that vary by deployment strategy. Using commercial APIs such as OpenAI’s GPT-4o mini, priced at \$0.3 per million input tokens and \$1.2 per million output tokens [^10] OpenAI API Pricing: https://openai.com/api/pricing/, a typical user interaction (200 input and 100 output tokens) costs approximately \$0.00018. If each of Netflix’s 301.6 million global subscribers received 50 recommendations daily, the resulting API cost would amount to approximately \$2.7 million per day. These costs must be weighed against potential revenue gains. In subscription-based models like Netflix, the average revenue per user in North America was about \$17.17 in 2024 (Singh 2025). In advertisement-driven platforms like YouTube, creators typically earn \$5 to \$7 per 1,000 ad views (Marshall 2024). Even modest improvements in user engagement could yield substantial returns, emphasizing the importance of careful cost-benefit analysis when scaling LLM integration.
Further, while LLMs present substantial opportunities to enhance recommendation quality and operational efficiency, successful deployment requires careful management of associated technical, economic, and ethical challenges. In particular, LLM-generated recommendations carry risks related to biased outputs (Feuerriegel et al. 2023; Bashardoust et al. 2024). In movie recommendation contexts, biases may manifest as an underrepresentation of specific genres, cultures, or demographic perspectives. For example, the corpus generated in our study tended to favor Western-centric narratives disproportionately. Additionally, perceived biases or cultural insensitivities arising from skewed training data or adversarial prompt manipulation can erode user trust and pose significant reputational risks to platforms.